A bullet filled with water has a mass of 1.3kg and is whirled vertically in a circle of radius 1.2m.The speed of the bucket at the bottom of the loop is 7.0ms–¹ . Clearly indicate on the bucket the direction of the force of

1.gravity(Fg)

2.tension(Ft)

Then calculate the acceleration


1
Expert's answer
2021-10-08T08:11:29-0400

"a=a_c=\omega^2r=\frac{v^2}r=40.8~\frac m{s^2},"

"F_g=mg=12.7~N,"

"F_t=m(g+a)=65.8~N."